# Analysis of the $q\bar{q}\to Z^{*}\to hA\to 4\tau$ process within the lepton-specific 2HDM at the LHC

### v2

@@ V Conclusions
In this paper, we have investigated the feasibility of hadro-producing the light Higgs scalar $h$ and pseudoscalar $A$ through $Z^{*}$ exchange (initiated by $q\bar{q}$ pairs) at the LHC in the 2HDM Type-X (or lepton specific). We have focused on the $4\tau$ final state, where two $\tau$’s decay leptonically to form a pair of SS leptons (electrons and/or muons) and two $\tau$’s decay hadronically (thus producing two jets). We have proven that this kind of signal could be found (or excluded) at the end of Run 3 of the LHC, following a dedicated selection based on both a kinematic and GBDT/TMVA analysis. In order to motivate future studies of this signature, we have also proposed 9 BPs and processed them through a realistic MC analysis for a centre-of-mass energy of $14$ TeV and (integrated) luminosity $L~{}=~{}\text{300}~{}\text{fb}^{-1}$, thus emulating actual LHC conditions.
@@ Acknowledgments
SM is supported in part through the NExT Institute and the STFC Consolidated Grant ST/X000583/1. SS is supported in full through the NExT Institute and acknowledges the use of the IRIDIS High Performance Computing Facility, and associated support services, at the University of Southampton, in the completion of this work. YW’s work is supported by the Natural Science Foundation of China Grant No. 12275143, and the Fundamental Research Funds for the Inner Mongolia Normal University Grant No. 2022JBQN080. QSY is supported by the Natural Science Foundation of China under the Grants No. 11875260 and No. 12275143.

### v3

@@ V Discussion and Conclusions
In this paper, we have investigated the feasibility of hadro-producing the light Higgs scalar $h$ and pseudoscalar $A$ through $Z^{*}$ exchange (initiated by $q\bar{q}$ pairs) at the LHC in the 2HDM Type-X (or lepton-specific). We have focused on the $4\tau$ final state, where two $\tau$’s decay leptonically to form a pair of SS leptons (electrons and/or muons), and two $\tau$’s decay hadronically (thus producing two jets). In order to motivate future studies of this signature, we have also proposed 9 BPs and processed them through a realistic MC analysis for a centre-of-mass energy of $14$ TeV and (integrated) luminosity $L=\text{300}~\text{fb}^{-1}$, thus emulating actual LHC conditions. We have proven that this kind of signal could be found (or excluded) at the end of Run 3 of the LHC, following a dedicated selection based on both a kinematic and GBDT/TMVA analysis.
Although our study primarily focused on LHC sensitivity, the large $\text{tan}\beta$ regions examined are also of interest for resolving the long-standing muon g2g-2 anomaly. BPs 5–9 with $\text{tan}\beta$ larger than 100 could resolve this anomaly by enhancing one-loop (${{H}}/{{A}}/{{H}}^{\pm}$) and two-loop (Barr-Zee) contributions. Future LHC data may narrow the viable explanations for the g2g-2 discrepancy. If the LHC excludes these high $\text{tan}\beta$ regions, it would also constrain the 2HDM Type-X explanation for g2g-2.
In the sign/background analysis, we have deliberately avoid using the tau taggings to reduce the fluctuation of our analysis. Here let’s address the effects of hadronic tau tagging briefly. Obviously, to introduce hadronic tau taggings would suppress both signal and background proportionally. For example, the sign mismeasured leptons from $Zjj$ cam be suppressed effectively, although such type of background can also be suppressed efficiently by the kinematic observables introduced in this work. Nevertheless, it is worthy of noting that after the pre-selection of same sign leptons, the dominant background is from $ZZ$ which can mimic our signal events mostly. Let’s assume a work point with a hadronic tau tagging efficiency $0.6$. Then, to impose 1 haodronic tau tagging can reduce the significance by a factor $0.77$, while impose two hadronic tau taggings can reduce the significance by a factor of $0.6$. As shown in Tab. VII, even after taking two hadronic tau tagging, the discovery potential of BPs introduced in this work can be achievable at the LHC when the luminosity is sufficient.
@@ Acknowledgments
SM is supported in part through the NExT Institute and the STFC Consolidated Grant ST/X000583/1. SS is supported in full through the NExT Institute and acknowledges the use of the IRIDIS High Performance Computing Facility, and associated support services, at the University of Southampton, in the completion of this work. YW’s work is supported by the Natural Science Foundation of China Grant No. 12275143, Central Guidance for Local Science and Technology Development Fund Project No.2025ZY0020 and the Fundamental Research Funds for the Inner Mongolia Normal University Grant No. 2022JBQN080. QSY is supported by the Natural Science Foundation of China under the Grants No. 11875260 and No. 12275143.